Whats the meaning of these lines:the angular momentum can have values such as h/2π,2h/2π,3h/2π….,etc.but it cannot have a fractional value.

Grade:12

7 Answers

Akash Kumar Dutta
10 years ago

net torque is zero during circular moiton due to presence of only central forces as a result angular momentum is conserved.
ang mom=(mv)r=pr=hr/(lambda)
now the wavelenth must be a part of the circumferance 2(pi)r/n
putiing the value.
ang mom=nh/2pi

adityanarayanan madhusudanan
10 years ago
According to Bohr''s second postulate, the electrons revolving in nuclear orbits of principal quantum numbers ''n'' can have only those values of angular momentum which are integral multiples of h/2p (i.e. mvr = nh/2p) . Therefore, fractional values of angular momentum are not allowed.  
Roshan Mohanty
10 years ago

We know that anugalar momentum can have value n*h/2*pie where n is a integer..so fractinaol values nt allowed

Prajwal kr
10 years ago

It means that angular momentum of an electron cannot just take any value. It is specific for different shells. It assume only dicreet values and thus the angular momentum of an electron is quantisized:

Angular momentum=mvr=nh/2pi

Akash Kumar Dutta
10 years ago

what havnt u undertood...??? u disapproved it!??
was it the wavelenght=2pie r/n ??
ok heres it :- the (wavelenght) .(a natural no.)=total circumference=2.pie.r
[this must be satisfied in order the electron''s orbit is completed]
hence,ang. mom.=nh/2pi
it depicts the ang mom. can only take integral values of h/2.pi
Plz approve

Abhishekh kumar sharma
10 years ago

According to Bohr''''s second postulate, the electrons revolving in nuclear orbits of principal quantum numbers ''''n'''' can have only those values of angular momentum which are integral multiples of h/2p (i.e. mvr = nh/2p) . Therefore, fractional values of angular momentum are not allowed.

pushpendra bansal
10 years ago

 angular momentum of electron in orbit is multiple of nH/2∏

where n is prinipal quantum number